You want to dilute a 6.00 M HCl solution to 3.01 M. You need 500. mL of the diluted solution. How much stock solution (6.00 M HCl) must you use? Give your answer in mL.
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 6 M

M2 = 3.01 M

V2 = 500 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (3.01*500)/6

V1 = 251 mL

Answer: 251 mL